OpenStudy (anonymous):

A polynomial function can be written as (x + 2)(x + 3)(x - 5). What are the x-intercepts of the graph of this function? (2, 0), (3, 0), (-5, 0) (-2, 0), (-3, 0), (5, 0) (2, 0), (3, 0), (5, 0) (-2, 0), (-3, 0), (-5, 0)

OpenStudy (lucaz):

x-intercepts are the roots, the values of x that makes y=0. the polynomial is already factored y=(x + 2)(x + 3)(x - 5). what are the falues of x that makes (x+2)=0, (x+3)=0 and (x-5)=0?

OpenStudy (anonymous):

idk this is all that the question asked i have no clue on how to do it

OpenStudy (lucaz):

(x+2)=0, when x is -2 you have: (-2+2) wich is zero

OpenStudy (lucaz):

so -2 is one of the roots, for (x+3) if x=-3 then (-3+3)=0, -3 is another root

OpenStudy (lucaz):

for (x-5) when x=5: (5-5)=0, the roots are -2, -3 and 5

OpenStudy (anonymous):

thanknu

OpenStudy (lucaz):

yw